Calculate the energy associated with 3rd orbital of H atom

Grade:11

1 Answers

Sudhir Bishnoi
6 years ago
E=-13.6(z^2÷n^2) eV
Here for 3rd orbital of H atom
        z=atomic number=1
        n=orbital number=3
Putting these values in eq
We get E=-13.6(1^2÷3^2)
             E=-13.6÷9
             E=1.51 eV 
             E=1.51*1.6*10-19
             E=2.41*10-19J